# Dynamics of Moral Behavior in Heterogeneous Populations of Learning Agents

### v6

@@ Appendix B Cooperation by Each Player Type across Populations
![[Uncaptioned image]](extracted/5942968/plots/appendix/manypopulaitons_actions_aggSelfish_avgacrossruns.png)
![[Uncaptioned image]](extracted/5942968/plots/appendix/manypopulaitons_actions_aggUtilitarian_avgacrossruns.png)
![[Uncaptioned image]](extracted/5942968/plots/appendix/manypopulaitons_actions_agganti-Utilitarian_avgacrossruns.png)
![[Uncaptioned image]](extracted/5942968/plots/appendix/manypopulaitons_actions_aggDeontological_avgacrossruns.png)
![[Uncaptioned image]](extracted/5942968/plots/appendix/manypopulaitons_actions_aggmalicious_Deontological_avgacrossruns.png)
@@
![[Uncaptioned image]](extracted/5942968/plots/appendix/manypopulaitons_actions_aggVirtueEthics_equality_avgacrossruns.png)
![[Uncaptioned image]](extracted/5942968/plots/appendix/manypopulaitons_actions_aggVirtueEthics_inequality_avgacrossruns.png)
![[Uncaptioned image]](extracted/5942968/plots/appendix/manypopulaitons_actions_aggVirtueEthics_kindness_avgacrossruns.png)
![[Uncaptioned image]](extracted/5942968/plots/appendix/manypopulaitons_actions_aggVirtueEthics_aggressiveness_avgacrossruns.png)

### v7

@@ Appendix B Cooperation by Each Player Type across Populations
![[Uncaptioned image]](extracted/6136686/plots/appendix/manypopulaitons_actions_aggSelfish_avgacrossruns.png)
![[Uncaptioned image]](extracted/6136686/plots/appendix/manypopulaitons_actions_aggUtilitarian_avgacrossruns.png)
![[Uncaptioned image]](extracted/6136686/plots/appendix/manypopulaitons_actions_agganti-Utilitarian_avgacrossruns.png)
![[Uncaptioned image]](extracted/6136686/plots/appendix/manypopulaitons_actions_aggDeontological_avgacrossruns.png)
![[Uncaptioned image]](extracted/6136686/plots/appendix/manypopulaitons_actions_aggmalicious_Deontological_avgacrossruns.png)
@@
![[Uncaptioned image]](extracted/6136686/plots/appendix/manypopulaitons_actions_aggVirtueEthics_equality_avgacrossruns.png)
![[Uncaptioned image]](extracted/6136686/plots/appendix/manypopulaitons_actions_aggVirtueEthics_inequality_avgacrossruns.png)
![[Uncaptioned image]](extracted/6136686/plots/appendix/manypopulaitons_actions_aggVirtueEthics_kindness_avgacrossruns.png)
![[Uncaptioned image]](extracted/6136686/plots/appendix/manypopulaitons_actions_aggVirtueEthics_aggressiveness_avgacrossruns.png)